# Correlations in Motion: A Simple Response-Based Analysis of Traffic Flow

### v3

@@ 1 Introduction
By conditioning responses on defined congestion indicators, we can extract how congestion affects downstream traffic, and do so in a way that reveals both propagation velocity and oscillation patterns, all from empirical time series data. While many established traffic models, such as fluid-dynamical theories [7, 9], gas-kinetic approaches [8, 11], and car-following theories [10, 12, 17], offer understanding of the mechanics of traffic flow, they often involve complex parameter estimation and limited generalizability. Similarly, cellular automata [13, 14, 15, 16] and three-phase traffic theory [18, 19, 20] have successfully reproduced emergent congestion patterns, but at the cost of considerable model complexity.
In contrast, response functions offer a complementary perspective: a straightforward way to empirically quantify how traffic reacts to prior disruptions. Recent studies have shown their potential in capturing temporal and spatial effects of congestion using velocity as the primary observable [21, 23, 25]. However, their broader capacity to quantify the interplay between flow, density, and velocity, and to systematically assess the spatiotemporal spread of congestion, remains largely unexplored. In this study, we address that by applying response functions to real motorway data from North Rhine–Westphalia, Germany, and investigate their ability to characterize traffic dynamics across multiple observables.
@@ 2.1 Data
[…]
@@ 2.2 Response functions and congestion indicators
$\varepsilon^{(0)}_{j}(t)=\begin{cases}1,&\text{if }0\leq v_{j}(t)\leq 20\ \text{km/h},\\ 0,&\text{else,}\end{cases},$ $\varepsilon^{(1)}_{j}(t)$ =$=\begin{cases}1,&\text{if }20\ \text{km/h}<v_{j}(t)\leq 40\ \text{km/h},\\ 0,&\text{else,}\end{cases}$ (6)
$\varepsilon^{(2)}_{j}(t)=\begin{cases}1,&\text{if }40\ \text{km/h}<v_{j}(t)\leq 60\ \text{km/h},\\ 0,&\text{else,}\end{cases},$ $\varepsilon^{(3)}_{j}(t)$ =$=\begin{cases}1,&\text{if }0\leq v_{j}(t)\leq 60\ \text{km/h},\\ 0,&\text{else.}\end{cases}$
@@ 4.2 On- and off-ramp propagation of congestion and generic response features
Notwithstanding the possibility to clearly identify congestion propagation through on- or off-ramps in both cases in Fig. 5, we find the unexpected result of the density response exhibiting a maximum for section 10 at interchange Herne. Lacking any obvious traffic-related reason or systematic error, the unexpected result may be caused by missing values in the analyzed data. For the indicator section 9 about 14% of that data are missing. For section 10 only 4% of the flow data and 9% of the velocity data are missing. The fluctuating courses of the flow responses also suggest a possible impact of low data quality on the anomalous results. Unfortunately, we cannot corroborate this point due to a lack of other available data, e.g. on construction sites. At this point, it is clear that data quality and detector availability limit our study. Both the characteristics of the minima (and maxima) and the fluctuations in the response functions depend directly on the data quality. In addition, a higher resolution of the data could provide more accurate results, but the measurements of the detectors take place over a time interval of one minute. For this reason, the detectors themselves limit our study.
@@ 4.4 The propagation velocity of congestion
[…]

### v4

@@ 1 Introduction
By conditioning responses on defined congestion indicators, we can extract how congestion affects downstream traffic, and do so in a way that reveals both propagation velocity and oscillation patterns, all from empirical time series data. While many established traffic models, such as fluid dynamical theories [7, 9], gas kinetic approaches [8, 11], and car following models [10, 12, 17], have provided deep insights into the mechanics of traffic flow and have been instrumental in understanding the conditions under which congestion arises, they often rely on specific model assumptions and parameter calibrations that may limit their generalizability. Cellular automata [13, 14, 15, 16] and three phase traffic theory [18, 19, 20] have successfully reproduced a wide range of emergent traffic patterns, including stop and go waves and capacity drops, often in high spatial resolution. However, these models can become complex when extended to real world networks or when applied to empirical validation.
Response functions offer a complementary perspective. Rather than aiming to model the full system from first principles, they empirically measure how traffic variables at one location respond to past disruptions elsewhere, directly from data. This allows for a transparent and model independent quantification of how congestion propagates in space and time. In this way, response functions do not replace established models but rather enhance our ability to test, interpret, and constrain them using observable system behavior. Especially for questions involving the temporal ordering and spatial spread of disturbances, response functions provide a robust, data driven framework that can be used in conjunction with more detailed simulation based approaches.
Recent studies have shown their potential in capturing temporal and spatial effects of congestion using velocity as the primary observable [21, 23, 25]. However, their broader capacity to quantify the interplay between flow, density, and velocity, and to systematically assess the spatiotemporal spread of congestion, remains largely unexplored. In this study, we address that by applying response functions to real motorway data from North Rhine–Westphalia, Germany, and investigate their ability to characterize traffic dynamics across multiple observables.
@@ 2.1 Data
is determined from the traffic flow $q_{sl}^{\text{(m)}}(t)$ and the velocity $v_{sl}^{\text{(m)}}(t)$ for cars and trucks with $\text{m}=\text{car},\text{tr}$. The sum of the respective quantities of cars and trucks provides the total traffic flow and density
@@ 2.2 Response functions and congestion indicators
[…]
@@ 4.2 On- and off-ramp propagation of congestion and generic response features
Notwithstanding the possibility to clearly identify congestion propagation through on- or off-ramps in both cases in Fig. 5, we find the unexpected result of the density response exhibiting a maximum for section 10 at interchange Herne. Lacking any obvious traffic-related reason or systematic error, the unexpected result may be caused by missing values in the analyzed data. For the indicator section 9 about 14% of that data are missing. For section 10 only 4% of the flow data and 9% of the velocity data are missing. The fluctuating courses of the flow responses also suggest a possible impact of low data quality on the anomalous results. Unfortunately, we cannot corroborate this point due to a lack of additional data, e.g. on construction sites or detector maintenance. At this point, it is clear that data quality and detector availability limit our study. Both the characteristics of the minima (and maxima) and the fluctuations in the response functions depend directly on the data quality. Due to occasional detector failures or temporary deactivations during construction, fully consistent data sets are difficult to obtain in practice. To mitigate this, we rely on averaging over a large number of independent days, which helps reduce the impact of isolated missing values and ensures stable response patterns. In addition, a higher resolution of the data could provide more accurate results, but the measurements of the detectors take place over a time interval of one minute. For this reason, the detectors themselves also limit our study.
@@ 4.4 The propagation velocity of congestion
where $x=v,q,\rho$ denotes the type of response and $\sigma=\text{min,max}$ the type of local extremum that has to be taken into account.